OpenStudy (anonymous):

Evaluate the expression described below if the number is –3. Five more than a number tripled.

OpenStudy (yamyam70):

x = 5 + x^3

OpenStudy (anonymous):

???

OpenStudy (john_es):

\[5+3x\Rightarrow5+3\cdot(-3)=-4\]

OpenStudy (anonymous):

-4 not an answer?

OpenStudy (john_es):

With the statement you give, the answer should be -4 (as far as I know a number tripled means 3x)

OpenStudy (john_es):

Which are the options they give you?

OpenStudy (anonymous):

A.1 b.-1/2 c.-12 d.7/12

OpenStudy (anonymous):

oh sry its right lol. was looking at the wrong problem lol

OpenStudy (john_es):

Ok, no problem

OpenStudy (anonymous):

Evaluate the expression for x = –3. x^2+4x+3?

OpenStudy (anonymous):

lol they gave me a butch of these

OpenStudy (anonymous):

would you mind?

OpenStudy (john_es):

Ok, it is easy, the only thing you have to do is replacing x for the number they give, \[(-3)^2+4(-3)+3=9-12+3=0\]

OpenStudy (anonymous):

Which of the following algebraic expressions represents the verbal expression given below? The opposite of a number, raised to the fourth power, is divided by five.

OpenStudy (anonymous):

oh why i do paste dat.

OpenStudy (anonymous):

let me try this lol

OpenStudy (john_es):

Ok

OpenStudy (anonymous):

-x^2/5?

OpenStudy (anonymous):

that right?

OpenStudy (john_es):

I would write the following, \[(-x)^4/5\]

OpenStudy (anonymous):

how you do the exponents like that?

OpenStudy (anonymous):

thats new lol

OpenStudy (john_es):

You can do that with the LaTeX environment or the Equation button. I use LaTeX so the sintaxis is, (-x)^4/5 But you need to put \ [ at the beginning and \ ] at the end of the expression but without spaces.

OpenStudy (anonymous):

gtg

OpenStudy (anonymous):

thanx again